OpenStudy (anonymous):

find the standard form equation of a hyperbola with vertices at (9,0),(-9,0) and end of the conjugate axis at (0,9)(0,-9)

OpenStudy (amistre64):

find your center; then its just finding some a,b that fits the conditions ...

OpenStudy (amistre64):

what do you have for a general hyperbola equation in your material?

OpenStudy (anonymous):

I don't have much of anything, he just gave us the paper to do and tried to explain it to us but none of us understood

OpenStudy (amistre64):

where to begin then ..... you need to know the basic setup for a hyperbola. its like an ellipse, but with subtraction.\[\frac{(x-x_c)^2}{a^2}-\frac{(y-y_c)^2}{b^2}=1\] but the terms can be swapped if the graph opens in the other direction

OpenStudy (amistre64):

the midpoint (xc,yc) of your vertexes define your center. can you tell me the center?

OpenStudy (anonymous):

what do you mean (xc,yc)

OpenStudy (amistre64):

the midpoint (xc,yc) of your vertexes ... the center is defined as a point with coordinants x and y .... i just used xc and yc to avoid confusion with some other point

OpenStudy (amistre64):